The following shared servers remain down:
Data Recovery process are still proceeding on these servers. Unfortunately we are unable to offer an ETA at this point. We hope to be able to advise an ETA as the day proceeds.
Network Routing, Co-location, VPS & VDS
Changes to the network infrastructure are now completed. Most of the routing issues have been resolved and any clients that remain affected will see their routing issued resolved today. This should resolve any further server connectivity issues.
While we continue to assess network vulnerabilities, some port lock-downs will continue to be in place. We would ask that those of you affected by these measures please remain patient whilst we assess the various parts of the network.
Further, we would advise clients that they should ensure that all administrative password to their servers should be changed.